OpenStudy (anonymous):

Solve: 89 + 2t = 10t + 25

OpenStudy (anonymous):

Solving for t?

OpenStudy (anonymous):

I can help.

OpenStudy (anonymous):

you need the answers

OpenStudy (anonymous):

First we want to isolate 't', that's what we are looking for.

OpenStudy (anonymous):

Nope :)

OpenStudy (anonymous):

k :)

OpenStudy (anonymous):

89 + 2t = 10t + 25 Subtract 25 from both sides 64+2t=10t subtract 2t 64=8t divide by 8 64/8=8t/8 t=8

OpenStudy (anonymous):

We can check to see if its right

OpenStudy (anonymous):

Substitute 8=t, into your equation.

OpenStudy (anonymous):

89 + 2t = 10t + 25 89+2(8)=10(8)+25 89+16=80+25 105=105. t=8. is true

OpenStudy (anonymous):

If you had multiple choives for 't', just use substitution to find 't'. Other than that, just perform operations to isolate your variable. :)